For an object projected from ground with speed u horizontal range is two times the maximum height attained by it. The horizontal range of object is

Grade:12th pass

1 Answers

Lakshmi Vinod
6 years ago
R=2H....(Given)We know thatR=4Hcot(theta)2H=4Hcot(theta)1=2cot(theta)1=2×[1/tan(theta)]Hence tan(theta)=2Now by using Range FormulaR=u×usin2(theta)/g=u×u/g×(2tan(theta)/1+tan×tan(theta)=u×u(2×2)/g(1+2×2)=4u×u/5g